City of Phoenix provides water to the Wilson's household. Last month, their water meter had a reading of 3,583 ccf. The previous month the meter reading was 3,576 ccf. City of Phoenix charges $8.93 per ccf for water. What were the Wilsons charged for water during this period?

1
Expert's answer
2021-09-21T00:14:57-0400

find the difference in reading of the meter for the two months=3583-3576=7ccf

Wilson's charges= 8.93"\times 7=\$62.51"